# Toward a Physical Theory of Intelligence

### v1

@@ Acknowledgements
The author acknowledges the use of OpenAI language models as a research and writing aid during the development, review, and refinement of the manuscript, and thanks the OpenAI team for developing these tools. All scientific claims, interpretations, and any errors remain the sole responsibility of the author. The author thanks his PhD supervisor and past mentors for discussions on topics related to artificial intelligence, dynamical systems, and information theory that helped shape the ideas in this work. This work was supported by a research studentship at the University of Edinburgh funded through the UK Research and Innovation (UKRI) Engineering and Physical Sciences Research Council (EPSRC).
Appendix A Example: Control–Induced Bit Flip
This appendix provides a minimal physical example illustrating the principles of Conservation-Congruent Encoding (CCE) and the physical origin of logical irreversibility. We consider a single logical bit realised by a controlled double–well potential whose metastable basins encode the logical states $L=\{0,1\}$. Logical transitions are implemented by deforming the potential through an external control parameter, allowing us to relate logical erasure to irreversible entropy flow in a conserved physical channel.
A.1 Potential and Control Schedule
Let $p\in P\subset\mathbb{R}$ denote the physical state and let $C_{t}\in C=\mathbb{R}$ be a control parameter that biases the energy landscape. The potential is
$U(p;C_{t})=ap^{4}-bp^{2}+c\,C_{t}\,p,$ (154)
with constants $a,b,c>0$. When $C_{t}=0$, the potential is symmetric and defines two metastable basins $(B_{0},B_{1})$ corresponding to the logical encodings under CCE. Negative $C_{t}$ favours basin $B_{0}$, while positive $C_{t}$ favours $B_{1}$.
A logical bit flip is implemented by a quasistatic control trajectory $C_{t}:[0,T]\to\mathbb{R}$ that smoothly tilts the potential from left to right:
$C_{t}=C_{\max}\,\sin\!\left(\frac{\pi t}{T}-\frac{\pi}{2}\right),$ (155)
so that $C_{0}=-C_{\max}$, $C_{T/2}=0$, and $C_{T}=+C_{\max}$. In the quasistatic limit $T\to\infty$, this protocol transports probability mass from $B_{0}$ to $B_{1}$ with arbitrarily small dissipation.
A.2 Physical Dynamics and Encoded Distributions
The microscopic dynamics of the system follow an overdamped Langevin equation
$\dot{p}_{t}=-\gamma^{-1}\,\partial_{p}U(p_{t};C_{t})+\sqrt{2D}\,\xi_{t},$ (156)
where $\gamma$ is a friction coefficient, $D=k_{B}T/\gamma$ is the diffusion constant, and $\xi_{t}$ is Gaussian white noise. The corresponding probability density $\nu_{t}(p)$ satisfies the Fokker–Planck equation
$\partial_{t}\nu_{t}(p)=\partial_{p}\!\Big(\gamma^{-1}(\partial_{p}U)\,\nu_{t}+D\,\partial_{p}\nu_{t}\Big).$ (157)
Initially, $\nu_{0}$ is concentrated in basin $B_{0}$, representing logical state “0”. As the control parameter $C_{t}$ is varied, the probability distribution $\nu_{t}$ continuously deforms and ultimately concentrates in basin $B_{1}$, representing logical state “1”. Under CCE, this process corresponds to a controlled deformation of metastable logical basins: no new encodings are introduced, and the logical transition is implemented entirely through changes in basin geometry.
A.3 Work, Entropy, and Irreversibility
The instantaneous work performed by the external control is
$\dot{W}(t)=\int\nu_{t}(p)\,\partial_{t}U(p;C_{t})\,dp=c\,\dot{C}_{t}\,\mathbb{E}_{\nu_{t}}[p].$ (158)
The total dissipated heat over the protocol is given by
$Q_{\mathrm{diss}}=\int_{0}^{T}\dot{W}(t)\,dt-\big(\mathbb{E}_{\nu_{T}}[U]-\mathbb{E}_{\nu_{0}}[U]\big),$ (159)
in accordance with the first law.
The coarse-grained physical entropy of the system is
$S_{\mathrm{phys}}(t)=-\,\alpha\int\nu_{t}(p)\,\ln\nu_{t}(p)\,dp,$ (160)
where $\alpha$ denotes the entropy scale associated with the conserved physical quantity through which irreversible information is exported (e.g. $\alpha=k_{B}$ for thermal channels). For a quasistatic transition between two equiprobable logical encodings, the coarse-grained entropy increases by
$\Delta S_{\mathrm{phys}}=\alpha\ln 2,$ (161)
reflecting the loss of one bit of distinguishability under the encoding map. This entropy increase must be exported through the conserved channel associated with the system’s irreversible dynamics. In thermal systems this reduces to the classical Landauer bound, but under Conservation–Congruent Encoding the result holds more generally for any conserved quantity supporting irreversible information loss.
A.4 Irreversibility and the CCE Interpretation
When the control protocol is not quasistatic, the metastable basins $B_{0}$ and $B_{1}$ may overlap transiently. During such intervals, the encoding map
$f:L\times C\;\to\;\mathcal{P}(P)$ (162)
is no longer invertible: distinct logical inputs induce physical distributions with overlapping support. This loss of distinguishability corresponds to logically irreversible processing.
Let $\dot{I}_{\mathrm{irr}}$ denote the rate at which such irreversible merges occur. Under CCE, each logical collapse of two equiprobable encodings is associated with a minimal entropy export of $\alpha\ln 2$. The minimal entropy production rate compatible with the observed loss of encodings is therefore
$\dot{S}_{\min}^{\mathrm{CCE}}=\alpha\ln 2\,\dot{I}_{\mathrm{irr}}.$ (163)
This simple double–well model thus makes explicit how logical irreversibility, physical irreversibility, and conserved–quantity dissipation are unified under CCE. Logical bit flips are thermodynamically irreversible unless performed in the ideal quasistatic limit; any finite–time protocol necessarily incurs positive entropy production proportional to the irreversibly erased information.
Appendix B Derivation of the Universal Trace Bound
This appendix provides a detailed derivation of the bound
$c_{t}\;\leq\;\tfrac{1}{2}\,\mathrm{tr}\!\big(\Sigma_{Z,t}\,\mathcal{F}^{(Z\!\to\!Y)}_{t}\big),\qquad\mathcal{C}_{T}\;\leq\;\frac{1}{T}\sum_{t=1}^{T}\tfrac{1}{2}\,\mathrm{tr}\!\big(\Sigma_{Z,t}\,\mathcal{F}^{(Z\!\to\!Y)}_{t}\big)$, (164)
which holds for all measurable channels $p(y_{t}\mid z_{t},\mathcal{H}_{t-1})$ that admit finite second moments and well-defined Fisher information matrices. The derivation requires no linearity, Gaussianity, or small-variance assumptions.
B.1 Setup
Fix a time index $t$ and condition on the history $\mathcal{H}_{t-1}=(Z^{t-1},Y^{t-1})$. Let $Z:=Z_{t}=\Gamma_{E}(E_{t})\in\mathscr{E}$ denote the environmental input port and $Y:=Y_{t}=\Gamma_{X}(X_{t})\in\mathscr{X}$ the corresponding output port. All expectations below are conditional on $\mathcal{H}_{t-1}$.
Define
$c_{t}$ $:=I(Z;Y\mid\mathcal{H}_{t-1})$
$=\mathbb{E}_{(Z,Y)\sim p(\cdot,\cdot\mid\mathcal{H}_{t-1})}\!\left[\log\frac{p(Z,Y\mid\mathcal{H}_{t-1})}{p(Z\mid\mathcal{H}_{t-1})\,p(Y\mid\mathcal{H}_{t-1})}\right]$
$=\mathbb{E}_{(Z,Y)\sim p(\cdot,\cdot\mid\mathcal{H}_{t-1})}\!\left[\log\frac{p(Y\mid Z,\mathcal{H}_{t-1})}{p(Y\mid\mathcal{H}_{t-1})}\right],$ (165)
$\Sigma_{Z}$ $=\mathrm{Cov}(Z\mid\mathcal{H}_{t-1}),$ (166)
$\mathcal{F}(z)$ $=\mathbb{E}_{Y\sim p(\cdot\mid z,\mathcal{H}_{t-1})}\!\left[\nabla_{z}\log p(Y\mid z,\mathcal{H}_{t-1})\nabla_{z}\log p(Y\mid z,\mathcal{H}_{t-1})^{\top}\right].$ (167)
We will show that $c_{t}$ is bounded by the trace inner product of the conditional input covariance $\Sigma_{Z}$ and a path-averaged Fisher matrix that reflects the channel’s local curvature with respect to $z$.
B.2 Step 1: Express $c_{t}$ as an expectation of KL divergences
For the conditional output distribution, define the random probability measure
$p_{Z}(y):=p(y\mid Z,\mathcal{H}_{t-1}),$ (168)
and define the corresponding mixture distribution
$p_{\mathrm{mix}}(y):=p(y\mid\mathcal{H}_{t-1}).$ (169)
By the definition of conditional mutual information,
$c_{t}=I(Z;Y\mid\mathcal{H}_{t-1})=\mathbb{E}_{Z\sim p(\cdot\mid\mathcal{H}_{t-1})}\!\left[D_{\mathrm{KL}}(p_{Z}\,\|\,p_{\mathrm{mix}})\right],$ (170)
where the expectation is taken over the law of $Z$ conditioned on $\mathcal{H}_{t-1}$.
B.3 Step 2: Replace the mixture by a pairwise comparison
The Kullback–Leibler divergence is convex in its second argument. Hence, for any random distribution $Q$ with $\mathbb{E}[Q]$ well defined,
$D_{\mathrm{KL}}\!\left(p_{Z}\,\middle\|\,\mathbb{E}[Q]\right)\;\leq\;\mathbb{E}\!\left[D_{\mathrm{KL}}(p_{Z}\,\|\,Q)\right].$ (171)
Let $\tilde{Z}$ denote an independent copy of $Z$ (under the conditional law), and take $Q:=p_{\tilde{Z}}$. Since $\mathbb{E}[p_{\tilde{Z}}]=p_{\mathrm{mix}}$, we obtain
[…]
Taking expectations over $Z$ yields
[…]
where $Z$ and $\tilde{Z}$ are independent draws (conditionally on $\mathcal{H}_{t-1}$).
B.4 Step 3: Information-geometric control of pairwise KL
For a smooth parametric family $\{p_{z}\}_{z\in\mathcal{Z}}$ with finite Fisher information, the following path-integral bound holds [1]:
$D_{\mathrm{KL}}(p_{z}\|p_{z^{\prime}})\;\leq\;\tfrac{1}{2}\,(z-z^{\prime})^{\top}\Bigg(\int_{0}^{1}\mathcal{F}(z_{s})\,ds\Bigg)(z-z^{\prime}),\qquad z_{s}:=z^{\prime}+s(z-z^{\prime}).$ (174)
This inequality follows by applying a second-order expansion of the KL divergence along the line segment $z_{s}$ and bounding the resulting curvature term by the Fisher information along the path.
Define the path-averaged Fisher matrix
$\overline{\mathcal{F}}(z,z^{\prime}):=\int_{0}^{1}\mathcal{F}(z_{s})\,ds\succeq 0.$ (175)
B.5 Step 4: Take expectations and separate spread from curvature
Substituting (174) into (173) gives
$c_{t}\;\leq\;\tfrac{1}{2}\,\mathbb{E}_{Z,Z^{\prime}}\!\Big[(Z-Z^{\prime})^{\top}\,\overline{\mathcal{F}}(Z,Z^{\prime})\,(Z-Z^{\prime})\Big].$ (176)
Using $\mathrm{tr}(aa^{\top}B)=a^{\top}Ba$, we can rewrite this as
$c_{t}\;\leq\;\tfrac{1}{2}\,\mathrm{tr}\!\Big(\mathbb{E}_{Z,Z^{\prime}}\!\big[(Z-Z^{\prime})(Z-Z^{\prime})^{\top}\,\overline{\mathcal{F}}(Z,Z^{\prime})\big]\Big).$ (177)
It is useful to isolate the input spread. Since $Z$ and $Z^{\prime}$ are i.i.d. (conditionally on $\mathcal{H}_{t-1}$),
$\mathbb{E}_{Z,Z^{\prime}}\!\big[(Z-Z^{\prime})(Z-Z^{\prime})^{\top}\big]=2\,\Sigma_{Z}.$ (178)
However, $\overline{\mathcal{F}}(Z,Z^{\prime})$ generally depends on $(Z,Z^{\prime})$, so the expectation in (177) does not in general factor into a product of expectations without additional assumptions. We therefore define the coupled curvature matrix
$\mathcal{G}_{t}\;:=\;\mathbb{E}_{Z,Z^{\prime}}\!\big[(Z-Z^{\prime})(Z-Z^{\prime})^{\top}\,\overline{\mathcal{F}}(Z,Z^{\prime})\big]\succeq 0,$ (179)
and obtain the universal trace bound
$c_{t}\;\leq\;\tfrac{1}{2}\,\mathrm{tr}(\mathcal{G}_{t}).$ (180)
B.6 Step 5: Average over time
Averaging the per-step bounds over the horizon $T$ yields
$\mathcal{C}_{T}[\bar{\mu}_{1:T}]\;\leq\;\frac{1}{T}\sum_{t=1}^{T}\tfrac{1}{2}\,\mathrm{tr}(\mathcal{G}_{t}),$ (181)
where $\mathcal{G}_{t}$ is the coupled spread–curvature matrix defined in (179).
B.7 Interpretation
  • The matrix $(Z_{t}-Z_{t}^{\prime})(Z_{t}-Z_{t}^{\prime})^{\top}$ captures the conditional spread of environmental inputs across the agent’s sensory port at time $t$.
  • The path-averaged Fisher matrix $\overline{\mathcal{F}}_{t}(Z_{t},Z_{t}^{\prime})$ quantifies the local curvature of the observation channel along the interpolation path between two input realizations $Z_{t}^{\prime}$ and $Z_{t}$, measuring sensitivity of the output distribution to changes in the input.
  • The coupled matrix $\mathcal{G}_{t}=\mathbb{E}[(Z_{t}-Z_{t}^{\prime})(Z_{t}-Z_{t}^{\prime})^{\top}\,\overline{\mathcal{F}}_{t}(Z_{t},Z_{t}^{\prime})]$ encodes the joint interaction between input variability and channel sensitivity. Its trace therefore bounds the instantaneous causal information throughput of the port at time $t$.
B.8 Assumptions
The bound (181) requires only the following regularity conditions:
  1. 1.
    For each $t$, the conditional likelihood $z\mapsto p(y_{t}\mid z,\mathcal{H}_{t-1})$ is twice differentiable, with finite Fisher information matrix $\mathcal{F}(z)$.
  2. 2.
    The conditional second moment $\mathbb{E}[\|Z_{t}\|^{2}\mid\mathcal{H}_{t-1}]$ is finite.
  3. 3.
    Expectations and derivatives may be interchanged (e.g. by dominated convergence).
No Gaussian, linear, or small-signal assumptions are required. For linear–Gaussian channels the bound is tight and reduces to the familiar quadratic Fisher-information form, while for general smooth stochastic channels it remains valid as a universal upper bound.
Appendix C Classical Thermodynamic Framing of Intelligence
This appendix provides a pedagogical special case of the general GENERIC formulation developed in the main text. We consider an isothermal system interacting with an environment at fixed temperature $T_{\mathrm{env}}$. The goal is to connect the reversible-dissipative decomposition of §7.1.2 to familiar entropy and free-energy balances in classical nonequilibrium thermodynamics.
C.1 Entropy and Free-Energy Balances
For a system with internal energy $U_{\mathrm{sys}}$, entropy $S_{\mathrm{sys}}$, and Helmholtz free energy $F_{\mathrm{sys}}=U_{\mathrm{sys}}-T_{\mathrm{env}}S_{\mathrm{sys}}$, the standard entropy balances for an open isothermal system are
$\dot{S}_{\mathrm{sys}}$ $=\dot{S}_{\mathrm{prod}}-\dot{S}_{\mathrm{export}},$ (182)
$\dot{S}_{\mathrm{env}}$ $=\dot{S}_{\mathrm{export}},$ (183)
$\dot{S}_{\mathrm{tot}}$ $=\dot{S}_{\mathrm{prod}}\geq 0,$ (184)
where $\dot{S}_{\mathrm{prod}}\geq 0$ is internal entropy production and $\dot{S}_{\mathrm{export}}$ is entropy flux to the environment.
Let $W_{T}$ denote useful work delivered by the system and $Q_{\mathrm{env}}$ the heat sent to the environment. The Clausius-Landauer relation gives
$Q_{\mathrm{env}}\;\geq\;k_{B}T_{\mathrm{env}}\,I_{\mathrm{irr}}-T_{\mathrm{env}}\,\Delta S_{\mathrm{sys}},$ (185)
where $I_{\mathrm{irr}}$ is the irreversible-information loss under Landauer-congruent encoding. Combining with $W_{T}=Q_{\mathrm{env}}-\Delta U_{\mathrm{sys}}$ and the definition of $F_{\mathrm{sys}}$ yields the instantaneous power bound
$\dot{W}(t)\;\leq\;k_{B}T_{\mathrm{env}}\,\dot{I}_{\mathrm{irr}}(t)-\dot{F}_{\mathrm{sys}}(t)-T_{\mathrm{env}}\,\dot{S}_{\mathrm{prod}}(t).$ (186)
This expression decomposes useful work into contributions from (i) irreversible information flow, (ii) changes in stored free energy, and (iii) internal entropy production.
C.2 Instantaneous Intelligence Bound
Define the instantaneous intelligence rate as the ratio of useful power to irreversible-information rate:
$\chi(t):=\frac{\dot{W}(t)}{\dot{I}_{\mathrm{irr}}(t)}.$ (187)
Normalising by $k_{B}T_{\mathrm{env}}$ and applying (186) gives
$\frac{\chi(t)}{k_{B}T_{\mathrm{env}}}\;\leq\;1-\frac{\dot{S}_{\mathrm{prod}}(t)}{\dot{I}_{\mathrm{irr}}(t)}-\frac{\dot{F}_{\mathrm{sys}}(t)}{k_{B}T_{\mathrm{env}}\,\dot{I}_{\mathrm{irr}}(t)}.$ (188)
The correction terms reflect (i) losses to internal dissipation and (ii) energy stored or released within the system. In the reversible limit where $\dot{S}_{\mathrm{prod}}\!\to\!0$ and $\dot{F}_{\mathrm{sys}}\!=\!0$, the classical Landauer bound $\chi(t)\!\to\!k_{B}T_{\mathrm{env}}$ is recovered.
C.3 Interpretation and Relation to the Metriplectic Framework
In the classical isothermal case, useful work increases as computations are routed through reversible pathways, and decreases when internal entropy is produced or free energy is accumulated rather than exported.
The metriplectic decomposition used throughout the main text provides a geometric generalisation of this picture. There, entropy production is generated exclusively by the dissipative flow $R\nabla\Phi$:
$\dot{S}_{\mathrm{prod}}(t)=\langle\nabla\Phi,R\nabla\Phi\rangle,\qquad\dot{I}_{\mathrm{irr}}(t)=\frac{\dot{S}_{\mathrm{prod}}(t)}{k_{B}},$ (189)
while reversible work is carried by the Hamiltonian flow $J\nabla H$. The reversible limit of the classical bound (188) corresponds precisely to the limit $R\nabla\Phi\to 0$ in Proposition 7.1.
Thus, the classical thermodynamic treatment arises as the special case of a more general reversible-dissipative structure: intelligent organisation corresponds to maximising useful work while minimising motion along the dissipative (entropy-producing) directions of the metriplectic dynamics.
Appendix D Oscillatory Bases, Fourier Analysis, and Reversible Computation
The metriplectic decomposition suggests a natural connection between reversible dynamics and classical Fourier analysis. Here we summarize this link and show how oscillatory modes provide a physically motivated basis for reversible predictive computation.
Fourier modes as reversible dynamical subsystems.
Consider a scalar stationary stochastic process $y(t)$ with finite second moments. Under standard regularity conditions, $y(t)$ admits a spectral representation of the form
$y(t)$ $=\int_{-\infty}^{\infty}\hat{y}(\omega)\,e^{i\omega t}\,d\mu(\omega),$ (190)
where $\hat{y}(\omega)$ are complex coefficients and $\mu$ is a spectral measure determined by the power spectral density $S_{y}(\omega)$. Each complex exponential $e^{i\omega t}$ can be expressed as a two-dimensional real oscillator,
$e^{i\omega t}$ $=\cos(\omega t)+i\sin(\omega t)=\begin{pmatrix}\cos(\omega t)\\ \sin(\omega t)\end{pmatrix}$. (191)
The corresponding dynamics satisfy
$\frac{d}{dt}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}=\omega\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}=J_{\omega}\nabla H(x),\qquad H(x)=\tfrac{1}{2}\|x\|^{2},$ (192)
where $J_{\omega}^{\top}=-J_{\omega}$ and no dissipative term is present. Eq. (192) is a special case of the reversible component $J\nabla H$: each Fourier mode evolves as an independent Hamiltonian oscillator executing phase rotation on a constant-energy circle.
In this representation, the environmental process $y(t)$ decomposes into a superposition of independent reversible oscillatory subsystems, each evolving under a norm-preserving flow generated by a skew-symmetric operator $J_{\omega}$. The Fourier basis therefore provides a canonical decomposition of environmental dynamics into elementary reversible components.
Reversible encoding as a unitary transform.
Suppose an agent maintains an internal state $x_{t}$ that encodes the recent history of $y(t)$ via a linear transform. A purely dissipative encoding (e.g. a leaky integrator) implements a non-unitary map that contracts volumes in state space and therefore erases information at each step. By contrast, encoding $y(t)$ in an oscillatory basis corresponds to a unitary (orthogonal) transform
$x_{t}$ $=U\,y_{t:t-\tau}$ (193)
for some windowed history $y_{t:t-\tau}$ and orthogonal matrix $U$ whose columns approximate Fourier modes. The subsequent time evolution of $x_{t}$ under the reversible dynamics $\dot{x}_{t}=J\nabla H(x_{t})$ preserves the norm and thus the mutual information between successive internal states:
$I[x_{t};x_{t+\Delta t}]_{\mathrm{rev}}$ $=I[x_{t};x_{t}],$ (194)
which equals $H(x_{t})$ in the noiseless limit. No information is erased by the reversible flow; the encoded spectrum is simply rotated in phase.
This corresponds to maximizing the reversible predictive information rate $\dot{I}_{\mathrm{pred,rev}}(J)$: once the relevant Fourier coefficients have been stored, the agent can propagate its internal model forward in time at negligible additional entropy cost, updating only when the external process deviates from the current prediction.
Dissipative filtering as repeated recomputation.
As a contrasting baseline, consider a scalar leaky integrator
$x_{t+1}$ $=(1-\alpha)x_{t}+\alpha y_{t},$ (195)
which implements a low-pass filter on $y_{t}$. Here the map $x_{t}\mapsto x_{t+1}$ is strictly contractive for $0<\alpha<1$, and the previous state is irreversibly overwritten by the new measurement. In information-theoretic terms, the agent partially discards its internal representation of the past at each step and reconstructs a new estimate from scratch, incurring a cost proportional to the information erased. The resulting irreversible information rate scales with the typical magnitude of the correction term $\Delta x_{t}^{\mathrm{irr}}=\alpha(y_{t}-x_{t})$.
By contrast, an oscillatory reservoir that encodes the dominant Fourier modes of $y(t)$ can achieve the same mean-square distortion with smaller dissipative corrections: most of the predictive structure is carried by the reversible phase rotation, and only phase or amplitude mismatches require irreversible updates. This realizes the inequality
$\dot{I}_{\mathrm{irr}}\;\geq\;R(D)-\dot{I}_{\mathrm{pred,rev}}(J)$ (196)
from Eq. (81): as the reversible basis captures more of the process spectrum, $\dot{I}_{\mathrm{pred,rev}}(J)$ increases and the irreversible rate required to maintain distortion $D$ decreases.
Summary.
Fourier analysis thus provides a concrete realization of the general principle developed in §7.1: oscillatory coupling is not merely a descriptive feature of biological dynamics, but an efficient basis for reversible predictive computation. Decomposing environmental signals into oscillatory modes allows an agent to (i) encode predictive structure via a unitary transformation, (ii) evolve this representation through reversible phase rotations with minimal entropy production, and (iii) confine irreversible computation to sparse corrective updates when predictions fail. This construction supplies an explicit spectral example of the energetic and information-theoretic arguments given in §7.1.2-7.1.4.

### v2

@@ Acknowledgements
The author acknowledges the use of OpenAI and Google language models, along with the Prism writing environment, as a research and writing aid during the development, review, and refinement of the manuscript. All scientific claims, interpretations, and any errors remain the sole responsibility of the author. The author thanks his PhD supervisor and past mentors for discussions on topics related to artificial intelligence, dynamical systems, and information theory that helped shape the ideas in this work. This work was supported by a research studentship at the University of Edinburgh funded through the UK Research and Innovation (UKRI) Engineering and Physical Sciences Research Council (EPSRC).
Appendix A Physical Realizations of Conservation-Congruent Encodings
The Conservation-Congruent Encoding (CCE) framework establishes that the irreversible collapse of a dynamically isolated equivalence class requires the export of a conserved quantity to the environment. The minimal generalised energetic cost of this operation is strictly bounded by the product of the exhaust channel’s characteristic fluctuation scale ($\alpha$) and its intensive conjugate force ($\mathcal{F}$), yielding the generalised Landauer bound:
$E_{\mathrm{min}}^{\mathrm{CCE}}=-(\mathcal{F}\alpha)\sum_{i=1}^{m}p_{i}\ln p_{i}.$ (193)
To demonstrate the absolute substrate neutrality of this framework, we detail how this single geometric constraint manifests across three distinct physical domains, carefully distinguishing the physical variables used to encode the information from the conserved quantities exported to erase it.
A.1 Electronic Systems (e.g., CMOS Memory)
In classical digital electronics, information is not encoded thermally; it is encoded electromagnetically. A logical state corresponds to the presence or absence of a macroscopic accumulation of electrical charge across a node capacitance.
However, the logical transitions between these charge states are dissipative by design. To irreversibly clear a register (e.g., discharging the capacitor to ground), the system must export energy into the surrounding phononic degrees of freedom (the thermal lattice).
  • Encoding Substrate: Macroscopic electrical charge ($q$) in a bistable voltage well.
  • Characteristic Scale ($\alpha$): The Boltzmann constant, $k_{B}$, defining the fundamental scale of the thermal fluctuations that the charge well must be robust against.
  • Conjugate Force ($\mathcal{F}$): The environmental temperature, $T$, of the thermal bath accepting the exhaust.
  • Physical CCE Cost: Erasing an equiprobable bit ($p_{0}=p_{1}=0.5$) forces the irreversible export of heat, perfectly recovering the classical Landauer limit:
    $E_{\mathrm{min}}=k_{B}T\ln 2.$ (194)
A.2 Chemical and Biological Systems (e.g., Molecular Gradients)
In biological computation, such as a synapse releasing neurotransmitters or a cell responding to an osmotic gradient, information is frequently encoded in the spatial compartmentalization of specific molecular species. The dynamically isolated equivalence classes are distinct concentration states.
When a cell irreversibly collapses this distinction—for example, by opening a membrane channel to allow the concentration gradient to equalise and “reset” the cellular memory—the relevant conserved quantity exchanged with the environment is particle number.
  • Encoding Substrate: Spatial distribution and concentration of a conserved chemical species (e.g., $\mathrm{Ca}^{2+}$ ions).
  • Characteristic Scale ($\alpha$): The scale is $1$ (for discrete single-particle analysis) or the ideal gas constant $R$ (if analysing macroscopic molar fluxes), defining the fundamental unit of the chemical channel.
  • Conjugate Force ($\mathcal{F}$): The chemical potential, $\mu$, which acts as the intensive driving force for particle exchange.
  • Physical CCE Cost: The minimal chemical work required to irreversibly merge two equiprobable compartmentalised states by exchanging particles with a reservoir is governed by the chemical potential:
    $W_{\mathrm{min}}=\mu\ln 2.$ (195)
A.3 Quantum Mechanical Systems (e.g., Spin Resonance)
In quantum information processing, an encoding may be physically instantiated in the intrinsic angular momentum of a particle (e.g., an electron spin oriented “up” or “down” along a quantization axis).
To irreversibly reset this spin to a known state (erasure), the system cannot simply dump heat; it must explicitly export angular momentum to an external field.
  • Encoding Substrate: The quantum spin state (angular momentum) of a particle.
  • Characteristic Scale ($\alpha$): The reduced Planck constant, $\hbar$, defining the fundamental scale of mechanical action and quantised angular momentum.
  • Conjugate Force ($\mathcal{F}$): The angular frequency of the applied external field, $\omega$ (such as the Larmor frequency in a magnetic resonance setup).
  • Physical CCE Cost: Irreversibly resetting the spin requires the extraction of angular momentum against the applied field, yielding:
    $E_{\mathrm{min}}=\hbar\omega\ln 2.$ (196)
In all three domains, the underlying physical mechanism remains identical: a macroscopic equivalence class is dynamically isolated to form a CCE, and its irreversible destruction demands a strict physical exhaust determined by the product of the channel’s specific conjugate force and characteristic scale ($\mathcal{F}\alpha$).